### v2

# xbx-supp-arxiv-1

Abstract
We introduce the XBX regression model, a continuous mixture of extended-support beta regressions for modelling bounded responses with boundary observations. The core building block of XBX regression is the extended-support beta distribution, a censored version of a four-parameter beta distribution with the same exceedance on the left and right of $(0,1)$. Hence, XBX regression is a direct extension of beta regression. We prove that beta regression and heteroscedastic normal regression with censoring at both 00 and $1$ — also known as the heteroscedastic two-limit tobit model in the econometrics literature — are special cases of extended-support beta regression, depending on whether a single extra parameter is zero or infinity, respectively. To overcome identifiability issues due to the similarity of the beta and normal distributions for certain parameter values, we shrink towards beta regression by letting the extra parameter have an exponential distribution with unknown mean. A Gauss-Laguerre quadrature approximation results in efficient likelihood-based estimation and inference procedures, which the betareg R package implements since version 3.2.0. We use XBX regression to analyze investment decisions in a behavioural economics experiment about the occurrence and extent of loss aversion. In contrast to standard approaches, we capture both the probability of rational behaviour and the mean of loss aversion. Extensive comparisons with alternative approaches illustrate the effectiveness of the new model.
Keywords: Boundary observations; approximate likelihood; tobit regression; loss aversion
1 Introduction
1.1 Bounded responses and beta regression
Bounded variables, such as rates, proportions, and fractional indices, are abundant in data-analytic applications. As with other variable types, linking them to available covariate information is a popular data-analytic task, for both inference on covariate effects and prediction. A range of statistical models has been developed in this direction, ranging from standard least squares approaches on transformed responses and other partially-specified models to fully parametric approaches; see Kieschnick and McCullough (2003) for an early review and discussion on available approaches.
From the available approaches, beta regression (Ferrari and Cribari-Neto, 2004; Smithson and Verkuilen, 2006) has emerged as the dominant likelihood-based approach for modelling bounded responses in terms of covariates. It has been successfully applied in diverse fields, including biology, ecology, economics and social sciences, finance, manufacturing, genetics, and engineering; see, for example, Allenbrand and Sherwood (2023), Leng et al. (2021), Namin et al. (2020), and Wang et al. (2015) for a few noteworthy applications, and Douma and Weedon (2019) and Geissinger et al. (2022) for recent articles making a case for its use in natural sciences and associated applied fields. The reason for its popularity is that beta regression respects the bounded range of the responses, and allows their mean and dispersion to be expressed in terms of covariates through separate regression structures. Such a construct offers high flexibility in the shape of the response distribution, capturing both the heteroscedasticity and the asymmetries observed in bounded responses.
Specifically, consider a random variable $Y$ with a beta distribution with density
$f_{\rm(B)}(y\mid\mu,\phi)=\frac{I(0<y<1)}{B(\mu\phi,(1-\mu)\phi)}y^{\mu\phi-1}(1-y)^{(1-\mu)\phi-1}$ (1)
where $\mu\in(0,1)$, $\phi>0$, $B(p,q)$ $(q,p>0)$ is the beta function (see, Abramowitz and Stegun, 1964, §6.2.1), and $I(A)$ is the indicator function taking value $1$ if $A$ is true, and zero, otherwise. In this parameterization, $\mathop{\rm E}(Y\mid\mu,\phi)=\mu$ and $\mathop{\rm var}(Y\mid\mu,\phi)=\mu(1-\mu)/(1+\phi)$ so that $\phi$ is a precision parameter. The beta regression model assumes that the observed responses $y_{1},\ldots,y_{n}$ are realizations of beta random variables $Y_{1},\ldots,Y_{n}$ that are independent conditionally on explanatory vectors $x_{1},\ldots,x_{n}$ and $z_{1},\ldots,z_{n}$, where $x_{i}$ and $z_{i}$ are $p$- and $q$-dimensional vectors, respectively, possibly sharing some or all components. As is apparent, we are assuming here that the observed responses take value between zero and one or have been transformed to do so as $y_{i}=(y_{i}^{\prime}-a)/(b-a)$, where $y_{i}^{\prime}$ are the original responses and $a$ and $b$ are the minimum and maximum value they can take. The mean and precision of $Y_{i}$ are then linked to the explanatory variables according to
$g_{1}(\mu_{i})=\eta_{i}=x_{i}^{\top}\beta\quad\text{and}\quad g_{2}(\phi_{i})=\zeta_{i}=z_{i}^{\top}\gamma\,,$ (2)
where $\beta=(\beta_{1},\ldots,\beta_{p})^{\top}$ and $\gamma=(\gamma_{1},\ldots,\gamma_{q})^{\top}$ are the vectors of regression parameters associated with the means and the precision of the beta variables, respectively.
The functions $g_{1}(\cdot)$ and $g_{2}(\cdot)$ are monotone functions, typically with the property of mapping the range of $\mu_{i}\in(0,1)$ and $\phi_{i}\in(0,\infty)$, respectively, to the real line. Usual choices for $g_{1}(\cdot)$ are the logit, probit, and generally any inverse of a cumulative distribution function, and for $g_{2}(\cdot)$ the log function.
The parameter vectors $\beta$ and $\gamma$ are typically estimated by maximum likelihood, and standard inferential procedures, such as likelihood ratio, Wald, and score tests, apply on the grounds of central limit theorems for the derivatives of the log-likelihood under assumptions on model adequacy. Grün et al. (2012) provide more details and describe a range of modelling strategies and methods based on the beta regression model.
1.2 Boundary observations
One important limitation of the beta regression model as defined by (1) and (2) is that likelihood-based estimation and inference is not possible when at least one of the observed responses is either zero or one. Specifically, depending on the parameter values, the density (1) at those observations is either zero or infinity. As a result and regardless of the sample size, the log-likelihood of the beta regression model is either zero or infinite or undetermined, and maximum likelihood estimation and the associated inferential procedures cannot be used. This is the reason that popular software for beta regression modelling, such as the betareg R package (Cribari-Neto and Zeileis, 2010; Grün et al., 2012) prior to version 3.2-0, returned an error when there is at least one response value that is zero or one.
One popular and straightforward approach to deal with boundary observations is by Smithson and Verkuilen (2006), who propose to transform the response to $(y_{i}+u)/(1+2u)$, with $u=1/\{2(n-1)\}$, and then fit the beta regression model in (1) and (2) with standard software. The transformed responses, then lie in $(0.5/n,1-0.5/n)\subset(0,1)$, and, as a result, there are no issues when fitting the beta regression model. The proposal in Smithson and Verkuilen (2006) has the advantage of preserving the parsimony and the convenience in estimation and inference that is provided by beta regression model in (1) and (2), but the value of $u$ is ad hoc and, as Example 1 demonstrates, can have a marked effect on inference.
Example 1:
Smithson and Verkuilen (2006, Example 3) models the contribution of nonverbal IQ and dyslexic versus nondyslexic status to the scores of 44 children on a test of reading accuracy. The scores were rescaled to lie in the unit interval, and 13 out of 44 children achieved a score of one. We transform the responses using $u=1/86$, and fit the beta regression model with $g_{1}(\mu_{i})=\log\{\mu_{i}/(1-\mu_{i})\}$, $g_{2}(\phi_{i})=\log(\phi_{i})$ in (2), where $\eta_{i}$ involves an intercept and the main and interaction effects of IQ and dyslexia status, and $\zeta_{i}$ involves only the main effects of IQ and dyslexia status. [^1] The numerical figures in Smithson and Verkuilen (2006, Table 5) can be reproduced by transforming only the boundary responses, and leaving the non-boundary ones unaltered. The same is true with Smithson and Merkle (2013, Table 6.2). We focus on the evidence against the omission of the interaction term.
Figure 1: Likelihood ratio and Wald statistics for the inclusion of an interaction term between dyslexia status and IQ in the setting of Example 1 versus $u$ ($\log_{10}$ scale). The vertical lines are the quantiles of the $\chi^{2}_{1}$ (likelihood ratio test) and normal distribution (two-sided Wald test) at levels $1.0\%$, $2.5\%$, $5.0\%$, $10.0\%$. The vertical dashed line is at $u=1/86$ as per the proposal in Smithson and Verkuilen (2006).
Figure 1 shows the value of the likelihood ratio and Wald statistics for $u$ ranging from $10^{-6}$ to $10^{-1}$. As is apparent, both tests are sensitive to the choice of $u$ with the evidence for the interaction term ranging from weak to strong. Notably, the closer the transformed responses get to the observed responses the weaker the evidence against the omission of an interaction term.
Another widely-used approach to deal with boundary observations is to construct two- or three-part hurdle models. Such models are formally mixtures of point masses at zero and/or one with a continuous distribution for the non-boundary observations. Hoff (2007), Cook et al. (2008), Ospina and Ferrari (2010), Calabrese (2012), Ospina and Ferrari (2012) define such a model by assuming a beta regression model for the non-boundary observations and use a regression structure for the hurdle part(s) of the model. The zoib (Liu and Kong, 2015) and gamlss (Rigby and Stasinopoulos, 2005) R packages can fit one- and two-part hurdle models [^2] Following the terminology set in the seminal work of Ospina and Ferrari (2010), the two- and three-part hurdle models are typically referred to as zero-or-one and zero-and-one inflated beta regression, respectively. We chose the term hurdle models instead, because the distributions in Ospina and Ferrari (2010) can be accurately described as zero and/or one being hurdles, which when passed a $(0,1)$ realization of a beta random variable is generated. In contrast, an inflated distribution is customarily understood as mixing the process that generates boundary values, with a distribution whose support already includes the boundary values (see, for example, Min and Agresti, 2002).. Hurdle models are a natural choice when it can be assumed that the boundary observations are generated by a distinct process from the one that generates the non-boundary observations. The separate regression structures for the hurdle part(s) of the model allow for linking the mixing probabilities directly with explanatory variables, offering both flexibility and convenience in estimation. If the beta regression part and the hurdle part(s) do not share any parameters then the likelihood is separable and can be maximized by separately estimating a binomial or multinomial regression model (depending on whether only one or both boundaries get point masses), and a beta regression model for the non-boundary observations.
Nevertheless, due to the many regression specifications involved in two- and three-part hurdle models, such models tend to be less parsimonious than the proposal in Smithson and Verkuilen (2006), which involves only mean and dispersion regression structures. Furthermore, hurdle models may be unnecessarily complex and hard to interpret, when there is merit in assuming that there is a single process generating the responses. For example, in Example 1, it is more natural to assume that there is a latent ability scale in terms of reading accuracy, which is linked to an observable test score with a known minimum and maximum value. A score of one represents that the child’s ability is sufficiently large to successfully complete all tasks in the given test. A hurdle model, in contrast, estimates separate processes for the perfect scores and all other scores (including those that are close to perfect).
Double censoring is another well-known approach for dealing with bounded-domain responses with boundary observations, particularly when it is natural to assume that there is a single process generating the response values. For double censoring, the observed bounded-domain response is assumed to be a realization of $Y=\max(\min(Y^{*},1),0)$, where $Y^{*}$ is a random variable with density function $f(\cdot\mid\theta)$. The marginal expectation of $Y$ is then a linear combination of the censoring probabilities and the expectation of the truncated distribution of $Y^{*}$ in $(0,1)$. Specifically,
$\mathop{\rm E}(Y)=P(Y^{*}\geq 1\mid\theta)+P(0<Y^{*}<1\mid\theta)\mathop{\rm E}(Y^{*}\mid 0<Y^{*}<1,\theta)\,.$ (3)
The likelihood about $\theta$ based on a sample $y_{1},\ldots,y_{n}$ is then
$\prod_{i:y_{i}=0}P(Y^{*}_{i}\leq 0\mid\theta)\prod_{i:y_{i}=1}P(Y^{*}_{i}\geq 1\mid\theta)\prod_{i:y_{i}\in(0,1)}f(y_{i}\mid\theta)\,,$ (4)
where $\prod_{A}a$ is $1$ if $A$ is the empty set. The standard approach towards defining a censored regression model is to put a regression structure on the components of $\theta$. The most widely-used censored regression model is the censored normal one, also known as the two-limit tobit. This model assumes that $Y^{*}_{1},\ldots,Y^{*}_{n}$ are independent conditionally on covariate vectors $x_{1},\ldots,x_{n}$, with $Y^{*}_{i}$ distributed according to a normal distribution with mean $x_{i}^{\top}\beta^{*}$ and variance $\sigma_{i}^{2}=\sigma^{2}$ (see, for example Rosett and Nelson, 1975; Maddala, 1983, Section 6.7, for definition and discussion). The heteroscedastic version of the two-limit tobit assumes that the variances are linked to explanatory variables as $g(\sigma_{i})=z_{i}^{\top}\gamma^{*}$, where $g(\cdot)$ is a monotone link function that maps $(0,\infty)$ to the real line (see, for example, Messner et al., 2016). Dixon and Sonka (1982) compare least squares and maximum likelihood estimation of the two-limit tobit in terms of estimates in an empirical application, and Hoff (2007) compares the two-limit tobit model to the semi-parametric approach in Papke and Wooldridge (1996) and the inflated beta distribution in Cook et al. (2008) for data envelopment analysis. Dixon and Sonka (1982), Hoff (2007), and Greene (2011, §19.3) touch on the interpretation of the parameters of the two-limit tobit. The disadvantage of the two-limit tobit is that it is less flexible than the beta distribution and cannot accommodate features that often occur in bounded-domain variables such as high skewness or kurtosis.
1.3 Our contribution
We develop the extended-support beta mixture (XBX) regression model, which accommodates boundary values of 00, $1$ or both 00 and $1$ for the responses, while preserving the parsimony and flexibility of beta regression. In contrast to the heteroscedastic two-limit tobit model, XBX regression has the flexibility of beta regression in the shape of the response distribution, allowing both for the heteroscedasticity and the asymmetries that may be observed in bounded responses.
The main building block for XBX regression is the extended-support beta (XB) distribution, which is constructed by first extending the support of the beta distribution to $(-u,1+u)$, $u>0$, and then censoring at both 00 and $1$. Thus, the XB distribution can be understood as a four-parameter beta distribution, restricted to symmetric exceedances and subsequently censored to obtain point masses on the boundaries of the unit interval. We show that the XB distribution converges to a beta distribution as $u\to 0$, and to a normal distribution censored in $[0,1]$, as $u\to\infty$. Hence, regression models based on the XB distribution bridge beta regression and censored normal models like the heteroscedastic two-limit tobit, at the expense of a single unknown constant $u$. Because the beta distribution can assume similar shapes to the normal distribution in its support for certain parameter values, it is natural to shrink XB towards beta. We achieve that through the XBX distribution, which is a continuous mixture of XB distributions with $u$ having an exponential distribution with unknown mean $\nu$. Then, as the mean $\nu$ of the exponential mixing distribution decreases, XBX gets closer to beta.
The likelihood of the XBX regression model can be conveniently approximated using a Gauss-Laguerre quadrature, giving rise to efficient estimation and inference procedures based on maximum approximate likelihood.
Finally, we present a case study highlighting the benefits of XBX regression. We revisit the analysis of an experiment from behavioural economics about the occurrence and extent of loss aversion (Glätzle-Rützler et al., 2015), which was originally analyzed using linear regression for the mean loss aversion only. Here, we demonstrate that the XBX model leads to qualitatively the same results for the effects of mean loss aversion, simultaneously allowing for economically relevant insights about the probability of rational behaviour.
The betareg R package, since version 3.2-0, provides methods for fitting and drawing inferences from XBX regression models.
Section 2 presents the XB distribution and shows that the beta and the censored normal distribution are its limiting cases. Section 3 introduces the XBX regression model, and discusses parameter estimation and inference using approximate likelihood. Section 4 illustrates the benefits of XBX regression through the analysis of a loss aversion experiment from behavioral economics. Section 5 presents a comprehensive simulation experiment comparing XBX regression to the heteroscedastic two-limit tobit model based on the continuous ranked probability score. We conclude with remarks and directions for further research in Section 6.
2 Extended-support beta distribution
[…]
$f_{\rm(B4)}(y\mid\mu,\phi,u_{1},u_{2})=f_{\rm(B)}\left(\frac{y-u_{1}}{u_{2}-u_{1}}\mid\mu,\phi\right)\frac{1}{u_{2}-u_{1}}\,,$ (5)
where ${u}_{1}<{u}_{2}$ are the boundary parameters.
[…]
Theorem 1:
Suppose that $Y$ is distributed according to the four-parameter beta distribution with density (5) and $u=-u_{1}=u_{2}-1$. Then,
  1. 1.
    If $u=0$, $Y$ has the beta density $f_{\rm(B)}(y\mid\mu,\phi)$ in (1).
  2. 2.
    If $\mu_{*}=\mathop{\rm E}(Y\mid\mu,\phi,u)\in\Re$ and $\sigma_{*}^{2}=\mathop{\rm var}(Y\mid\mu,\phi,u)\in(0,\infty)$ are fixed, then, as $u\to\infty$, $Y$ converges in distribution to $\mu_{*}+\sigma_{*}Z$, where $Z$ is a standard normal random variable.
Proof.
Part 1 is direct by setting $u_{1}=0$ and $u_{2}=1$ in (5). For part 2, let $\alpha=\mu\phi$ and $\beta=(1-\mu)\phi$. For fixed $\mu_{*}$ and $\sigma_{*}$, $\alpha$ and $\beta$ grow quadratically with $u$, so $\alpha\to\infty$ and $\beta\to\infty$ as $u\to\infty$. From the properties of the four-parameter beta distribution $\mu_{*}=(1+2u)\mu-u$ and $\sigma_{*}^{2}=(1+2u)^{2}\sigma^{2}$, where $\sigma^{2}=\mu(1-\mu)/(1+\phi)$. Hence, $(y+u)/(1+2u)=\mu+\sigma t$, where $t=(y-\mu_{*})/\sigma_{*}$. Then, writing the density (5) in terms of $t$ and multiplying it by $dy/dt$ gives that the density of $T=(Y-\mu_{*})/\sigma_{*}$ is
$h(t\mid\mu,\phi,u)=\frac{\sigma(\mu+\sigma t)^{\alpha-1}(1-\mu-\sigma t)^{\beta-1}}{B(\alpha,\beta)}I\left(\frac{-u-\mu_{*}}{\sigma_{*}}<t<\frac{1+u-\mu_{*}}{\sigma_{*}}\right)\,.$ (6)
For $u\to\infty$, the indicator function on the right hand side of (6) has value $1$ for any finite $t$. Furthermore, Moscovich et al. (2016, Lemma A.1) shows that the first factor on the right hand side of (6) converges to $e^{-t^{2}/2}/\sqrt{2\pi}$ for any $t$ as $\alpha\to\infty$ and $\beta\to\infty$. By Scheffé’s theorem, $T$ converges in distribution to $Z$, and by the continuous mapping theorem $Y$ converges in distribution to $\mu_{*}+\sigma_{*}Z$. ∎
[…]
$f_{\rm(XB)}(y\mid\mu,\phi,u)=\left\{\begin{array}[]{ll}\displaystyle F_{\rm(B)}\left(\frac{u}{1+2u}\mid\mu,\phi\right)\,,&\text{if }y=0\\[10.00002pt] \displaystyle f_{\rm(B)}\left(\frac{y+u}{1+2u}\mid\mu,\phi\right)\frac{1}{1+2u}\,,&\text{if }y\in(0,1)\\[10.00002pt] \displaystyle 1-F_{\rm(B)}\left(\frac{1+u}{1+2u}\mid\mu,\phi\right)\,,&\text{if }y=1\\ \end{array}\right.\,,$ (7)
where $F_{\rm(B)}(y\mid\mu,\phi)$ is the distribution function of the beta distribution with density (1). For the parameter settings considered in Figure 2, the density for the continuous part of the XB distribution is exactly equal to the dark grey regions in $(0,1)$, while the discrete part will consist of point masses at 00 and $1$ that are equal to the probability at the left and right of $(0,1)$.
A direct corollary of Theorem 1 is that the XB distribution coincides with the beta distribution for $u=0$, and that it converges to the normal distribution censored in $[0,1]$ as $u\to\infty$. Hence, regression models based on the XB distribution allow to parameterize the transition from beta regression to the heteroscedastic two-limit tobit, at the expense of a single extra parameter $u$. The beta distribution and the normal distribution limits of the XB distribution can assume similar shapes in $(0,1)$ for certain parameter settings. Hence, it is natural to consider shrinking $u$ to 00, or, equivalently, shrinking the XB distributions towards the beta distribution, to avoid any identifiability issues that may arise in the estimation of the parameters of the XB distribution, while preserving the flexibility offered by the beta distribution.
3 Continuous mixtures of extended-support beta distributions
3.1 Exponential extended-support beta mixture
Shrinkage of the XB distribution towards beta can be induced in a tuning-parameter-free way by constructing a continuous mixture of XB distributions. In this direction, we consider a continuous mixture of XB distributions, where $u$ has an exponential distribution with unknown mean $\nu$, that is $u\sim{\rm Exp}(\nu)$ and $Y\mid u\sim{\rm XB}(\mu,\phi,u)$. We denote the marginal distribution of $Y$ as ${\rm XBX}(\mu,\phi,\nu)$, which has density
[…]
where $f_{\rm(XB)}(y\mid\mu,\phi,u)$ is defined in (7). By its definition, for fixed $\phi$ and $\nu$, the XBX density is symmetric in that $f_{\rm(XBX)}(y\mid\mu,\phi,\nu)=f_{\rm(XBX)}(1-y\mid 1-\mu,\phi,\nu)$. Density (8) is not available in closed form, but it can be accurately and efficiently approximated using a Gauss-Laguerre quadrature rule (see, Abramowitz and Stegun, 1964, Equation 25.4.45) as
$f_{\rm(XBX)}(y\mid\mu,\phi,\nu)\simeq\sum_{t=1}^{T}W_{t}f_{\rm(XB)}(y\mid\mu,\phi,\nu Q_{t})\,,$ (9)
where $W_{t}$ and $Q_{t}$ $(t=1,\ldots,T)$ are weights and nodes whose calculation and derivation are through Laguerre polynomials.
Figure 3: Densities of the beta and XBX distribution for all combinations of $\mu\in\{0.5,0.7,0.9\}$, $\phi\in\{0.5,2,20\}$, and $\nu\in\{0.01,0.1,1\}$. The horizontal dotted line is at 1. The height of the vertical black segments represents the point mass at 00 and $1$ for the corresponding parameter setting.
Figure 3 shows the density in $(0,1)$ and the point masses at 00 and $1$ of the XBX distribution for a range of values for $(\mu,\phi,\nu)$. As expected, as $\nu\to 0$, the XBX distribution converges to the beta distribution. On the other hand, as $\nu$ increases the point masses at the boundaries increase, and the relative size of that increase at 00 and $1$ depends on the skewness of the XBX distribution.
3.2 XBX regression
The operations of mixing and censoring that are involved in the definition of the XBX distribution are interchangeable. Specifically, the XBX distribution can be equivalently defined as the distribution of $Y=\max(\min(Y^{*},1),0)$, where $Y^{*}\mid u\sim{\rm B4}(\mu,\phi,-u,1+u)$ , and $u\sim{\rm Exp}(\nu)$. The expectation and variance of $Y^{*}\mid u$ are $\mathop{\rm E}(Y^{*}\mid\mu,\phi,u)=(1+2u)\mu-u$ and $\mathop{\rm var}(Y^{*}\mid\mu,\phi,u)=(1+2u)^{2}\mu(1-\mu)/(1+\phi)$, respectively. The marginal expectation and variance of $Y^{*}$, when $u\sim{\rm Exp}(\nu)$, are directly obtained by the laws of total expectation and variance as
$\mathop{\rm E}(Y^{*}\mid\mu,\phi,\nu)$ $=(1+2\nu)\mu-\nu\,,$ (10)
$\mathop{\rm var}(Y^{*}\mid\mu,\phi,\nu)$ $=(2\mu-1)^{2}\nu^{2}+(1+4\nu+8\nu^{2})\frac{\mu(1-\mu)}{1+\phi}\,,$
respectively. Similarly to the heteroscedastic two-limit tobit, we can assume that $Y^{*}$ is a latent variable in $\Re$, which is linked to an observable response in $[0,1]$, where the boundary values of 00 and $1$ represent too small and too large values of the latent variable. As discussed in Section 1.2, this is a natural assumption in assessment scores and testing settings, like that of Example 1.
From expressions (10), the parameters $\mu$ and $\phi$ have direct interpretations in terms of location and precision of the distribution of $Y^{*}$, as they do in beta regression. In particular, the expectation of $Y^{*}$ is a linear function of $\mu$, with slope $(1+2\nu)$ and intercept $-\nu$. On the other hand, the variance of $Y^{*}$ decreases as $\phi$ increases, with the rate of decrease increasing quadratically with $\nu$.
So, an XBX regression with mean and dispersion effects for the latent distribution can directly be defined by linking $\mu$ and $\phi$ to covariates as in (2). From Figure 3, it is directly apparent that the response distribution in XBX regression can have much more flexible shapes than the censored normal distribution underlying the heteroscedastic two-limit tobit; see, also Figure LABEL:fig-supp:xbx-vs-cn and Figure LABEL:fig-supp:xbx-vs-cn-hist in Section LABEL:sec-supp:xbx-vs-cn of the supplementary material document for a side-by-side comparison of the XBX distribution to the doubly-censored normal distribution.
The expectation and variance of a random variable with XBX distribution are, of course, nonlinear functions of $\mu$, $\phi$ and $\nu$ as they depend on the probabilities of $Y^{*}$ being greater or equal to $1$ and $Y^{*}$ being in $(0,1)$, and on moments of the truncated distribution of $Y^{*}$ in $(0,1)$. Figure LABEL:fig-supp:xbx-e and Figure LABEL:fig-supp:xbx-v of Section LABEL:sec:xbx-e-v the supplementary material show the expectation and variance of the XBX distribution as a function of $\mu$ for a range of values for $\phi$ and $\nu$. For reference, Figure LABEL:fig-supp:xbx-e and Figure LABEL:fig-supp:xbx-v also show the expectation ($\mu$) and variance ($\mu(1-\mu)/\phi$) of the beta distribution as functions of $\mu$ for a range of values of $\phi$.
As expected, for small $\nu$, the expectation and variance of the XBX distribution are almost identical to those of the beta distribution. Notably, for all values of $\nu$ and $\phi$, the expectation of the XBX distribution is found to be a monotone increasing function of $\mu$ and rotationally symmetric at $(0.5,0.5)$. Hence, any interpretations of the signs of the coefficients in the regression structure for $\mu$ in XBX regression are as in standard beta regression, with the usual caveat that changes in covariates in the regression structure on $\mu$ impact also the variance of the response distributions. Like for the beta distribution, the variance of the XBX distribution as a function of $\mu$ is bounded from above for any $\phi$ and $\nu$, with the difference that it can assume bimodal shapes. Furthermore, as for beta, the variance of the XBX distribution decreases as $\phi$ increases. In contrast to beta, however, the mean of the XBX distribution increases with $\phi$ if $\mu>0.5$ and decreases if $\mu<0.5$, most notably for large values of $\nu$. So, in XBX regression, if a regression structure on $\phi$ is present alongside a regression structure on $\mu$, the associated covariates would also contribute to the mean of the response distribution. The same holds for other censored regression models, like the two-limit tobit, where regression structures on the variance of the latent variable, contribute to the mean of the response distribution.
3.3 Maximum approximate likelihood
From equation (9), the log-likelihood function of the XBX regression model can be approximated using the Gauss-Laguerre quadrature rule. The parameters are the regression coefficients $\beta$ and $\gamma$ from (2) as well as the log exceedance parameter $\xi=\log\nu$, to assure that all parameters have unrestricted ranges. This yields the approximate log-likelihood
$\ell^{\rm(a)}(\beta,\gamma,\xi)=\sum_{i=1}^{n}\log\sum_{t=1}^{T}W_{t}L_{i}(\beta,\gamma,e^{\xi}Q_{t})\,,$ (11)
where $L_{i}(\beta,\gamma,u)=f_{\rm(XB)}(y_{i}\mid\mu_{i},\phi_{i},u)$ is the contribution of the $i$th observation to the likelihood of the corresponding regression model based on the XB distribution; see (7). The gradient of (11) is
$s^{\rm(a)}(\beta,\gamma,\xi)=\sum_{i=1}^{n}\frac{\sum_{t=1}^{T}W_{t}L_{i}(\beta,\gamma,e^{\xi}Q_{t})\nabla\log L_{i}(\beta,\gamma,e^{\xi}Q_{t})}{\sum_{t=1}^{T}W_{t}L_{i}(\beta,\gamma,e^{\xi}Q_{t})}\,.$ (12)
Implementing the approximate log-likelihood (11) for XBX regression is straightforward. The key ingredients are a routine to compute the weights and nodes of the Gauss-Laguerre quadrature, such as the gauss.quad() function of the statmod R package (Smyth, 2005), and routines to compute the density and distribution functions of beta random variables, such as the dbeta() and pbeta() R functions. From (12), the gradient of the approximate log-likelihood for XBX regression is also readily available, in light of expressions for the contributions $\nabla\log L_{i}(\beta,\gamma,\nu Q_{t})$ to the log-likelihood gradient of the corresponding regression model based on the XB distribution. A tedious but straightforward calculation gives that the gradient of $\log L_{i}(\beta,\gamma,u)$ with respect to $\beta$ is
$\nabla_{\beta}\log L_{i}(\beta,\gamma,u)=\left\{\begin{array}[]{ll}\displaystyle x_{i}^{\top}\phi_{i}d_{1,i}\frac{Q_{0,i}-R_{0,i}}{F_{0,i}}\,,&\text{if }y=0\\[10.00002pt] \displaystyle x_{i}^{\top}\phi_{i}d_{1,i}(\bar{T}_{i}-\bar{U}_{i})\,,&\text{if}y\in(0,1)\\[10.00002pt] \displaystyle x_{i}^{\top}\phi_{i}d_{1,i}\frac{R_{1,i}-Q_{1,i}}{1-F_{1,i}}\,,&\text{if }y=1\end{array}\right.\,,$ (13)
the gradient with respect to $\gamma$ is
$\nabla_{\gamma}\log L_{i}(\beta,\gamma,u)=\left\{\begin{array}[]{ll}\displaystyle z_{i}^{\top}d_{2,i}\frac{\mu_{i}Q_{0,i}+(1-\mu_{i})R_{0,i}}{F_{0,i}}\,,&\text{if }y=0\\[10.00002pt] \displaystyle z_{i}^{\top}d_{2,i}\{\mu_{i}(\bar{T}_{i}-\bar{U}_{i})+\bar{U}_{i}\}\,,&\text{if }y\in(0,1)\\[10.00002pt] \displaystyle z_{i}^{\top}d_{2,i}\frac{\mu_{i}Q_{1,i}+(1-\mu_{i})R_{1,i}}{1-F_{1,i}}\,,&\text{if }y=1\end{array}\right.\,,$ (14)
and that the partial derivative with respect to $u$ is
$\frac{\partial}{\partial u}\log L_{i}(\beta,\gamma,u)=\left\{\begin{array}[]{ll}\displaystyle\frac{f_{0,i}}{(1+2u)^{2}F_{0,i}}\,,&\text{if }y=0\\[10.00002pt] \displaystyle\frac{\mu_{i}\phi_{i}-1}{y_{i}+u}+\frac{(1-\mu_{i})\phi_{i}-1}{1-y_{i}+u}-\frac{2(\phi_{i}-1)}{1+2u}\,,&\text{if }y\in(0,1)\\[10.00002pt] \displaystyle\frac{f_{1,i}}{(1+2u)^{2}(1-F_{1,i})}\,,&\text{if }y=1\end{array}\right.\,.$ (15)
In expressions (13), (14), and (15), $d_{1,i}=d\mu_{i}/d\eta_{i}$, $d_{2,i}=d\phi_{i}/d\zeta_{i}$, and
$f_{j,i}=f_{(b)}(z(j,u)\mid\mu_{i},\phi_{i})\quad\text{and}\quad F_{j,i}=F_{(b)}(z(j,u)\mid\mu_{i},\phi_{i})\,,$
with $z(j,u)=(j+u)/(1+2u)$. In addition,
$\bar{T}_{i}=\log Y_{i}-\psi(\phi\mu_{i})+\psi(\phi_{i})\quad\text{and}\quad\bar{U}_{i}=\log(1-Y_{i})-\psi(\phi(1-\mu_{i}))+\psi(\phi_{i})\,,$
$Q_{j,i}=q(z(j,u)\mid\mu_{i},\phi_{i})$, and $R_{j,i}=R(z(j,u)\mid\mu_{i},\phi_{i})$ with
$q(z\mid\mu,\phi)$ $=F_{(B)}(z\mid\mu,\phi)\left\{\psi(\phi)-\psi(\phi\mu)+\log z\right\}-h(z\mid\mu,\phi)\,,$
$r(z\mid\mu,\phi)$ $=F_{(B)}(1-z\mid 1-\mu,\phi)\left\{\psi(\phi(1-\mu))-\psi(\phi)-\log(1-z)\right\}-h(1-z\mid 1-\mu,\phi)\,,$
and
$h(z,\mu,\phi)=\frac{z^{\mu\phi}}{\mu^{2}\phi^{2}B(\mu\phi,(1-\mu)\phi)}{{}_{3}}F_{2}(\mu\phi,\mu\phi,1-(1-\mu)\phi\,;\,\mu\phi+1,\mu\phi+1\,;\,z)\,,$
where $\psi(\cdot)$ is the digamma function, and ${}_{3}F_{2}(\cdot,\cdot,\cdot\,;\,\cdot,\cdot)$ is the generalized hypergeometric function.
Maximum approximate likelihood estimates can be computed using either gradient-free optimization algorithms or optimization algorithms that operate with numerical or analytical gradient approximations. Estimated standard errors can be computed as the diagonal of the inverse of the negative hessian of the approximate log-likelihood, which is the observed information matrix, and can be accurately approximated using numerical differentiation procedures on (11) or (12).
3.4 Inference
Hypothesis tests and confidence sets can be constructed using standard tools for likelihood inference. For example, letting $\theta=(\beta^{\top},\gamma^{\top})^{\top}$ and $I^{(a)}(\hat{\theta})$ be the negative hessian of the approximate log-likelihood (11), the hypothesis $R\theta=b$ for a known $c\times(p+q)$ martix $R$ can be tested by comparing the approximate likelihood ratio statistic
$2\left\{\ell^{(a)}(\hat{\beta},\hat{\gamma},\hat{\xi})-\max_{R\theta=b}\ell^{(a)}(\beta,\gamma,\xi)\right\}\,,$
or the Wald statistic
$(R\hat{\theta}-b)^{\top}\{RI^{(a)}(\hat{\theta})^{-1}R^{\top}\}^{-1}(R\hat{\theta}-b)\,,$
to the quantiles of their asymptotic distribution, that is $\chi^{2}$ with $c$ degrees of freedom.
4 Application: Loss aversion in investment decisions
4.1 Motivation
To illustrate the benefits of XBX regression models, we revisit the analysis of a behavioral economics experiment conducted and published by Glätzle-Rützler et al. (2015). The outcome variable is the proportion of tokens invested by high-school students in a risky lottery with positive expected payouts. Glätzle-Rützler et al. (2015) focused on the effects of several experimental factors on the mean investments, which reflect the players’ willingness to take risks. The latter study employed linear regression models, estimated by ordinary least squares (OLS) with standard errors adjusted for potential clustering and heteroscedasticity.
Here, we extend the analysis from Glätzle-Rützler et al. (2015) by employing a similar model for the mean investments but exploring distributional specifications that allow for a probabilistic, rather than mean-only, interpretation of the effects. From an economic perspective this is of interest because it allows to interpret both the mean willingness to take risks in this experiment, and the probability to behave like a rational Homo oeconomicus, who would invest (almost) all tokens in this lottery because it has positive expected payouts.
4.2 Data
The experiment was designed by Matthias Sutter and Daniela Glätzle-Rützler at Universität Innsbruck to study a phenomenon known as myopic loss aversion (Benartzi and Thaler, 1995). The classical setup for demonstrating this is a lottery with positive expected payouts where individuals typically invest less than the maximum possible (loss aversion), and where investments further decrease if the individuals can make (many) short-term investment decisions (myopia or short-sightedness) as opposed to (fewer) long-term decisions.
The experiment involved high-school students from Schwaz and Innsbruck in Tyrol, Austria, who participated as individuals or as teams of two students. The subjects, consisting of 385 individuals and 185 teams of two, could invest $X\in\{0,\dots,100\}$ tokens in each of nine rounds in a lottery. The payouts were $2100+5X/$ 2 tokens with probability $1/3$ and $100-X$ tokens with probability $2/3$. Thus, the expected payouts were $6100+X/$ 6 tokens, and, as a result, a Homo oeconomicus would always invest the maximum of $X=100$ tokens in each round. It was observed that 1.5% of the subjects never invested anything, and 5.3% always invested everything.
Interest is in linking the proportion $Y_{i}\in[0,1]$ of the total tokens invested by the $i$th subject in the nine rounds out of the $900$ available tokens to explanatory information. The explanatory information we consider here consists of the subject characteristics and their interactions that Glätzle-Rützler et al. (2015) found to be relevant in the experiment. These are Grade ($G_{i}$ with value 0 and 1, according to whether the the $i$th subject is from lower grades 6–8 or upper grades 10–12, respectively), Arrangement ($T_{i}$, with values 0 and 1, according to whether the $i$th subject is an individual or a team of two, respectively), Age ($A_{i}$ the age of the player or the average age of the team of players in years), Sex ($S_{i}$, with value 1 if the player was male, or the team included at least one male student, and 0, otherwise), and the interactions of Grade with Arrangement and Age.
4.3 Models
We compare four different models for the proportion of the total tokens invested. All models assume that $Y_{1},\ldots,Y_{n}$ $(n=570)$ are independent conditionally on the covariate information, and that the mean parameter for the distribution of $Y_{i}$ is linked to a linear predictor of the form
$\eta_{i}=\beta_{1}+\beta_{2}G_{i}+\beta_{3}T_{i}+\beta_{4}A_{i}+\beta_{5}S_{i}+\beta_{5}G_{i}T_{i}+\beta_{7}G_{i}A_{i}\,.$ (16)
The first model is a normal linear regression model (N) under which $Y_{i}$ has a normal distribution with mean $\mu_{i}=\eta_{i}$ and variance $\sigma_{i}=\sigma^{2}$. This model corresponds to the OLS approach used by Glätzle-Rützler et al. (2015). The second model we consider is the heteroscedastic two-limit tobit model, which is based on the censored normal distribution (CN) and described in Section 1.2, and where the latent variable $Y_{i}^{*}$ has mean $\mu_{i}=\eta_{i}$ and variance $\sigma_{i}$ satisfying $\log(\sigma_{i})=\zeta_{i}$ with
$\zeta_{i}=\gamma_{1}+\gamma_{2}G_{i}+\gamma_{3}T_{i}+\gamma_{4}S_{i}\,.$ (17)
For simplicity, the regression structure (17) involves only main effects, excluding $A_{i}$ because $G_{i}$ already captures the main age differences between the lower and upper grades. The third model we consider is the beta regression model (B) after adjusting the responses using the ad-hoc scaling of Smithson and Verkuilen (2006) (as described in Section 1.2) with mean $\mu_{i}$ satisfying $\log\{\mu_{i}/(1-\mu_{i})\}=\eta_{i}$, with $\eta_{i}$ as in (16), and precision parameter $\phi_{i}$ satisfying $\log\phi_{i}=\zeta_{i}$, with $\zeta_{i}$ as in (17). Finally, we also consider the extended-support beta mixture model (XBX) where the mean and precision parameters $\mu_{i}$ and $\phi_{i}$ (see (10)), respectively, have the same specification as in the beta regression model.
4.4 Estimates
Coefficient estimates (and estimated standard errors) along with log-likelihood, Akaike and Bayes information criteria (AIC and BIC) for all four models are provided in Table 1 and Table 2, respectively. Separate tables are used because the coefficients from N and CN use an identity link for the mean parameter whereas B and XBX use a logit link. In addition, the log-likelihood, and, hence, AIC and BIC, are comparable only between CN and XBX because those two models have the same support for the response variable, that is the unit interval with point masses at 0 and 1.
Table 1: Coefficient estimates, estimated standard errors (in parentheses), maximized log-likelihood, AIC, and BIC for the normal model (N) and the censored normal model (CN) of Section 4.3.
Covariate N CN
$\mu_{i}$ $\sigma_{i}$ $\mu_{i}$ $\log(\sigma_{i})$
(Intercept) $0.284$ $0.246$ $0.250$ $-1.467$
(0.157) (0.154) (0.047)
$G_{i}$ (Grade) $-0.844$ $-0.982$ $0.167$
(0.281) (0.317) (0.066)
$T_{i}$ (Arrangement) $0.063$ $0.065$ $-0.138$
(0.030) (0.029) (0.072)
$A_{i}$ (Age) $0.012$ $0.014$
(0.012) (0.012)
$S_{i}$ (Sex) $0.104$ $0.113$ $0.212$
(0.023) (0.025) (0.068)
$G_{i}T_{i}$ (Grade $\times$ Arrangement) $0.151$ $0.167$
(0.045) (0.050)
$G_{i}A_{i}$ (Grade $\times$ Age) $0.046$ $0.054$
(0.019) (0.020)
Log-likelihood $-8.8$ $-75.5$
AIC $33.6$ $173.0$
BIC $68.3$ $220.8$
The coefficient estimates (and the estimated standard errors) in the $\mu_{i}$ regression structure for the N and CN models are similar, with the ones for N being slightly smaller than those for CN. This resembles the attenuation bias phenomenon when estimating regression parameters using least squares in the presence of bounded responses with boundary observations (see, for example, Winkelmann and Boes, 2006, Section 7.2).
Table 2: Coefficient estimates, estimated standard errors (in parentheses), maximized log-likelihood, AIC, and BIC for the beta model (B) and the extended beta mixture model (XBX) of Section 4.3.
Covariate B XBX
$\text{logit}(\mu_{i})$ $\log(\phi_{i})$ $\text{logit}(\mu_{i})$ $\log(\phi_{i})$ $\log(\nu)$
(Intercept) $-1.414$ $1.194$ $-0.865$ $1.756$ $-2.273$
(0.620) (0.084) (0.519) (0.128) (0.245)
$G_{i}$ (Grade) $-2.943$ $-0.553$ $-3.096$ $-0.316$
(1.252) (0.104) (1.053) (0.131)
$T_{i}$ (Arrangement) $0.225$ $0.406$ $0.208$ $0.325$
(0.117) (0.122) (0.099) (0.145)
$A_{i}$ (Age) $0.091$ $0.049$
(0.049) (0.041)
$S_{i}$ (Sex) $0.455$ $-0.555$ $0.379$ $-0.484$
(0.099) (0.113) (0.084) (0.136)
$G_{i}T_{i}$ (Grade $\times$ Arrangement) $0.655$ $0.567$
(0.200) (0.169)
$G_{i}A_{i}$ (Grade $\times$ Age) $0.151$ $0.169$
(0.081) (0.068)
Log-likelihood $94.4$ $-71.8$
AIC $-166.8$ $167.7$
BIC $-119.0$ $219.8$
Similarly, the coefficient estimates in the $\log\{\mu_{i}/(1-\mu_{i})\}$ regression structure for the B and XBX are comparable and lead to similar qualitative interpretations. The estimates for XBX (and their estimated standard errors) are slightly smaller than for B. The reason is that the estimated effects in B are slightly exaggerated to push fitted values towards the boundaries, while the extra parameter $\nu$, estimated to be $0.1$, in XBX offers more flexibility in capturing the boundary observations.
Figure 4: Hanging rootograms for the fitted models in Table 1 and Table 2. The rootograms compare the histogram estimate of the marginal distribution of the proportions of tokens invested with the corresponding expected frequencies from the fitted models. The square root of the expected frequencies are shown as dots connected by a solid line (in black) and the square root of the observed frequencies are hanging from the points as grey bars. The quality of the fit is determined by checking whether the bars hang down to the horizontal line at zero or whether they are clearly too large or too small. The dashed horizontal lines are the Tukey warning limits at ±$\pm 1$ (see Kleiber and Zeileis, 2016, for details).
4.5 Assessment of fitted models
The differences between the models become more apparent when considering the fitted distributions. Figure 4 shows hanging rootograms (see Kleiber and Zeileis, 2016, for details) for the fitted models in Table 1 and Table 2. The rootograms compare the empirical marginal distributions of the proportion of tokens invested to the implied distributions from the fitted models. The quality of the fit is determined by checking whether the bars hang down to the horizontal line at zero or whether they are clearly too large or too small. As is apparent, models N and B fit poorly in the tails. For N the support of the fitted distributions is the real line and, as a result, the expected frequencies in intervals close to 00 or $1$ are larger and smaller, respectively, than what is observed. The corresponding expected frequencies for B are larger than what is observed, while the expected frequencies at intervals around 0.5 are too small compared to what is observed. In contrast, models CN and XBX fit very well with all but the second bar from the left hanging close to the zero reference line and within the Tukey limits. The misfit of the second bar is larger for CN compared to XBX, which is also reflected by both AIC and BIC being smaller for XBX; see Table 1 and Table 2.
Testing whether the removal of dispersion effects $\log(\phi)=\zeta_{i}$ in model XBX, with $\zeta_{i}$ as in (17), is worthwhile returns an approximate likelihood ratio statistic of $20.614$ and a Wald statistic of $21.251$ on $3$ degrees of freedom, which is strong evidence against the hypothesis that $\log(\phi)=\gamma_{1}$. We get the same qualitative conclusions from model CN about the removal of variance effects $\log(\sigma)=\zeta_{i}$, with a likelihood ratio statistic of $19.620$ and a Wald statistic of $19.243$ on $3$ degrees of freedom.
4.6 Effects
Models XBX and CN allow more economically relevant interpretations of probability effects than the N and B models. For illustration, we focus on the team arrangement effect for a subsample with a large share of very rational subjects: male players or teams with at least one male, in grades 10–12, and between 15 and 17 years of age. About $52\%$ of the subjects in that subsample invested more than half of their total tokens.
Table 3: Estimated arrangement effects for subjects that are male or teams with at least one male, in grade 10–12 and 16 years of age. The rows for $E(Y)$ show model-based estimates of the expected proportion of tokens invested, and the rows for $P(Y>0.95)$ show model-based estimates of the probability to behave very rationally by investing more than $95\%$ of the tokens. All effects are contrasted to the corresponding empirical quantities obtained from the subsample of subjects that are male or teams with at least one male, in grades 10–12, and between 15 and 17 years of age. The rows for the parameters show the estimates of the distributional parameters for each arrangement setting.
Arrangement Empirical N CN B XBX
$E(Y)$ Individual $0.482$ $0.461$ $0.470$ $0.492$ $0.471$
Team $0.665$ $0.675$ $0.674$ $0.700$ $0.686$
$P(Y>0.95)$ Individual $0.087$ $0.023$ $0.075$ $0.125$ $0.072$
Team $0.207$ $0.131$ $0.194$ $0.249$ $0.185$
Parameters Individual $\mu$ $0.461$ $0.465$ $\mu$ $0.492$ $0.475$
$\sigma$ $0.246$ $0.337$ $\phi$ $1.091$ $2.602$
$\nu$ $0.103$
Team $\mu$ $0.675$ $0.696$ $\mu$ $0.700$ $0.662$
$\sigma$ $0.246$ $0.294$ $\phi$ $1.637$ $3.602$
$\nu$ $0.103$
Figure 5: Estimated arrangement effects in Table 3.
Table 3 shows the model-based estimates of the expected proportion of tokens invested ($E(Y)$) and the probability to behave very rationally by investing more than $95\%$ ($P(Y>0.95)$) of the tokens, contrasted to the corresponding empirical quantities from the subsample with age between 15 and 17 years. The mean and probability effects are also visualized in Figure 5; see, also, Figure LABEL:fig:lossaversion-cdf in the supplementary material document for the fitted cumulative distribution functions. All models do a reasonable job in estimating $E(Y)$, with the largest deviations observed for model B. Nevertheless, the censored models XBX and CN estimate the probabilities to behave very rationally much better than B and N do. Model N, due to its unlimited support, underestimates the variance and, hence, underestimates that probability, while model B with limited support and no point masses overestimates that probability. As is apparent using a model with a good fit is essential for estimating probability effects, especially so at the tails of bounded domains. Table LABEL:tab:lossaversion-left-prob in the supplementary material document shows the model-based estimates for $P(Y<0.05)$, which, albeit not being as important as $P(Y>0.95)$ in terms of its economic interpretation, further illustrates the quality of the estimated probability effects from XBX and CN at the tails.
4.7 Three-part hurdle models
In our analysis, we have not considered the three-part hurdle model of Ospina and Ferrari (2010) (“zero-and-one-inflated” beta) for two reasons: (a) from an economic point of view, it is not plausible that the effects driving “fully rational” behavior ($Y=1$) are different from the effects driving “very rational” behavior (say, $Y<0.95$ or $Y=1$); (b) the hurdle needs substantially more parameters than XBX due to the regression structures for modelling the probability of boundary and non-boundary observations, whose estimation may also not be reliable. For example, for the loss aversion data, employing a multinomial logistic regression model for the probabilities of boundary and non-boundary observations with the same covariates as those in the linear predictor (16) or (17), and model B for the non-boundary observations, results in a three-part hurdle model with 25 or 19 parameters, respectively, instead of the 12 parameters that XBX requires. Notably, maximum likelihood estimation results in infinite estimates for the parameters of either multinomial logistic regression model (see Table LABEL:tab:lossaversion-beta01-1 and Table LABEL:tab:lossaversion-beta01-2 in the supplementary material document for estimates and estimated standard errors) due to data separation. Also, the BIC values from both three-part hurdle model are larger than those of the CN and XBX models.
5 XBX versus heteroscedastic two-limit tobit regression
In the application of Section 4, XBX regression resulted in a slightly better fit than the CN model. In this section, we carry out a numerical study to compare the performance of XBX and CN regression in a more systematic way.
Theorem 1 shows that regression models based on the XB distribution allow to parameterize the transition from the beta distribution to the censored normal distribution with a single parameter $u$. On the other hand, XBX regression is a continuous mixture of XB distributions that shrinks towards a beta regression model. Hence, a comparison of XBX regression with heteroscedastic two-limit tobit regression can be based on data sets from the XB distributions. For a given $(\beta_{1},\beta_{2},\gamma_{1},\gamma_{2},u)^{\top}$, we let $\log\{\mu_{i}/(1-\mu_{i})\}=\beta_{1}+\beta_{2}x_{i}$ and $\log\phi_{i}=\gamma_{1}+\gamma_{2}x_{i}$, where $x_{i}=2(i-1)/(n-1)-1\in[-1,1]$ $(i=1,\ldots,n)$, and simulate response values $y_{1},\ldots,y_{n},y_{1}^{\prime},\ldots,y_{n}^{\prime}$ as realizations of independent random variables $Y_{1},\ldots,Y_{n},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$ with $Y_{i}\sim{\rm XB}(\mu_{i},\phi_{i},u)$, and $Y_{i}^{\prime}\sim{\rm XB}(\mu_{i},\phi_{i},u)$. We then fit the XBX and CN models on $y_{1},\ldots,y_{n}$ with covariates $x_{1},\ldots,x_{n}$ in both mean and variance and precision regression specifications. The predictive performance of each model is evaluated as
$S=\sum_{i=1}^{n}C(\hat{F}_{i},y_{i}^{\prime})\,,$ (18)
[…]
Table 4: Intervals $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$ considered in the simulation study of Section 5, and the implied values of $\beta_{1}$, $\beta_{2}$, $\gamma_{1}$ and $\gamma_{2}$ (in three decimal places).
$\mu_{1}$ $\mu_{n}$ $\beta_{1}$ $\beta_{2}$
$0.05$ $0.25$ $-2.022$ $0.923$
$0.05$ $0.50$ $-1.472$ $1.472$
$0.05$ $0.75$ $-0.923$ $2.022$
$0.05$ $0.95$ $-0.000$ $2.944$
$0.25$ $0.50$ $-0.549$ $0.549$
$0.25$ $0.75$ $0.000$ $1.099$
$0.25$ $0.95$ $0.923$ $2.022$
$0.50$ $0.75$ $0.549$ $0.549$
$0.50$ $0.95$ $1.472$ $1.472$
$0.75$ $0.95$ $2.022$ $0.923$
$\phi_{1}$ $\phi_{n}$ $\gamma_{1}$ $\gamma_{2}$
$0.5$ $10.0$ $0.805$ $1.498$
$0.5$ $20.0$ $1.151$ $1.844$
$0.5$ $50.0$ $1.609$ $2.303$
$5.0$ $100.0$ $3.107$ $1.498$
$10.0$ $20.0$ $2.649$ $0.347$
$20.0$ $50.0$ $3.454$ $0.458$
$50.0$ $100.0$ $4.259$ $0.347$
Because $x_{1},\ldots,x_{n}$ define an equispaced grid of size $n$ on $[-1,1]$, the parameters $\beta_{1},\beta_{2},\gamma_{1},\gamma_{2}$ are uniquely determined by $\mu_{1}$, $\mu_{n}$, $\phi_{1}$, and $\phi_{n}$. Also, for $\mu_{1}\leq\mu_{n}$ and $\phi_{1}\leq\phi_{n}$, it necessarily holds that $\mu_{i}\in[\mu_{1},\mu_{n}]$ and $\phi_{i}\in[\phi_{1},\phi_{n}]$. Table 4 shows the intervals $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$ we consider in the simulation experiment, along with the implied parameter values. The intervals cover narrow and wide ranges for $\mu$, being either close to the boundary values or around 0.5, and narrow and wide intervals for $\phi_{i}$, with both low and high values of the precision parameters.
We set $n=500$. In order to avoid trivial response configurations, from the possible combinations of $u\in\{2^{-6},2^{-5},\ldots,2\}$, and $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$ in Table 4, we consider those for which the average probability of non-boundary response values in $(0,1)$ is more than $0.05$. We then simulate $100$ independent vectors $(y_{1},\ldots,y_{n},y_{1}^{\prime},\ldots,y_{n}^{\prime})$ per parameter setting, and compute the relative change in $S$, $S_{\rm CN}/S_{\rm XBX}-1$, when moving from the XBX to the CN model, based on their maximum likelihood fits. Positive values of the relative change correspond to $S_{\rm CN}>S_{\rm XBX}$, and, thus, worse predictive performance of the CN model relative to the XBX regression model.
Figure 6 shows the average relative change for a subset of the parameter settings we consider; see, Figure LABEL:fig-supp:xbx-vs-htobit in the supplementary material document, for all settings. The XBX model is performing the same or outperforms the CN model in the majority of settings. CN is found to perform slightly better when the response distributions have small variance (large values of $\phi$) for moderate values of $u$. Those settings produce datasets where the relationship between the non-boundary responses and covariates is well-modelled as linear; see Figure LABEL:fig-supp:xb-data-1 to Figure LABEL:fig-supp:xb-data-4 in the supplementary material document, which show simulated data sets for all parameter settings. The two models start performing similarly again for large $u$, when the response values start concentrating on the boundaries.
Figure 6: Average relative change in $S$ in (18) ($S_{\rm CN}/S_{\rm XBX}-1$) when moving from XBX to CN, based on $100$ samples of size $n=500$ for each combination of $u\in\{2^{-6},2^{-5},\ldots,2\}$, and a subset of $[\mu_{1},\mu_{n}]$ (columns) and $[\phi_{1},\phi_{n}]$ (rows) intervals in Table 4. Positive values of the relative change correspond to $S_{\rm CN}>S_{\rm XBX}$, and, thus, worse predictive performance of the CN model relative to the XBX regression model. The dashed vertical lines are the smallest values of $u$ for which the average probability of boundary observations exceeds $0.95$ for each combination of $[\mu_{1},\mu_{n}]$ and $[\phi_{1},\phi_{n}]$.
The experimental setting in the current section has been used to compare beta regression on rescaled responses according to the proposal in Smithson and Verkuilen (2006) and the normal linear regression model. Figure LABEL:fig-supp:xbx-vs-beta and Figure LABEL:fig-supp:xbx-vs-ols in the Section LABEL:sec:XBX-vs-beta-ols of the supplementary material document show the average relative change in $S$ when moving from XBX regression to each of those models as a function of $u$ for all parameter settings in Table 4. In both cases XBX regression performs best and increasingly better as $u$ increases.
6 Concluding remarks
We have introduced XBX regression for modelling bounded responses with or without boundary observations. At the core of the XBX regression model is the XB distribution, which is a censored version of a four-parameter beta distribution with the same exceedance $u$ on the left and right of $(0,1)$. We prove that both beta regression with dispersion effects and the heteroscedastic two-limit tobit model are special cases of XB regression, depending on whether $u$ is zero or infinity, respectively. To avoid identifiability issues that may arise in estimating $u$ due to the similarity of the beta and normal distribution for specific parameter settings that result in low skewness and kurtosis, we define XBX regression as a shrunken version of XB regression towards beta regression, by letting $u$ have an exponential distribution with mean $\nu$. The associated marginal likelihood can be conveniently and accurately approximated using a Gauss-Laguerre quadrature rule, resulting in efficient estimation and inference procedures. In Section 5, we produce evidence that XBX regression can substantially outperform heteroscedastic two-limit tobit regression when the common features of skewness and kurtosis of bounded-domain variables are encountered, and perform only slightly worse when skewness and kurtosis are small.
In Section 4, we use XBX regression to analyze investment decisions in the behavioral economics experiment of Glätzle-Rützler et al. (2015), where the occurrence and extent of loss aversion are of interest. In contrast to standard approaches, XBX regression can simultaneously capture the probability of rational behavior and the mean amount of loss aversion.
The XBX regression model can be used to model bounded responses with or without boundary observations in arbitrary intervals $[a,b]$, for known $a$ and $b$ by simply transforming the observed responses $y_{i}^{\prime}$ to $y_{i}=(y_{i}-a)/(b-a)$. Furthermore, the assumption of an exponential distribution for $u$ has been made just for convenience. The exponential distribution has only one parameter, and its simple form allows the use of Gauss-Laguerre quadrature for maximum approximate likelihood, also giving rise to simple marginal moments for the latent uncensored variable; see (10). Other assumptions for the distribution of $u$ can also result in shrinkage towards beta regression. We believe that the benefits in applications will, at best, be marginal, at the expense of potentially complicating estimation and key characteristics of the latent distribution. It is also tempting to allow for regression structures in $\nu$. However, by (10), $\nu$ controls the rate that the marginal variance of the latent distribution changes with the precision parameter $\phi$. As a result, regression structures on both $\nu$ and $\phi$ can offer little extra flexibility compared to a regression structure on $\phi$ and estimation of $\nu$. Additional flexibility can be achieved in a more interpretable manner by exploring finite mixtures of XBX regression models, similarly to how Grün et al. (2012) defined finite mixtures of beta regression models.
The betareg R package, since version 3.2-0, provides methods for fitting and drawing inference from XBX regression models. Bayesian inference for XBX regression is straightforward through template modelling frameworks such as JAGS (Plummer, 2003) and Stan (Carpenter et al., 2017), due to the alternative definition of XBX regression in Section 3 as the censored version of a mixture of four-parameter beta regressions with the same exponentially-distributed exceedance on the left and right of $(0,1)$. Methods based on Stan are provided by the brms R (Bürkner, 2017) package since version 2.23.
Supplementary material
The supplementary material is available at https://github.com/ikosmidis/XBX-supplementary and provides the supplementary material document referenced above, and instructions and scripts to reproduce all numerical results and figures in the main text and the supplementary material document.
Declarations
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.